# Entropy-based random models for hypergraphs

### v2

@@ Abstract
Network theory has primarily focused on pairwise relationships, disregarding many-body interactions: neglecting them, however, can lead to misleading representations of complex systems. Hypergraphs represent an increasingly popular alternative for describing polyadic interactions: our innovation lies in leveraging the representation of hypergraphs based on the incidence matrix for extending the entropy-based framework to higher-order structures. In analogy with the Exponential Random Graphs, we name the members of this novel class of models Exponential Random Hypergraphs. Here, we focus on two explicit examples, i.e. the generalisations of the Erdös-Rényi Model and of the Configuration Model. After discussing their asymptotic properties, we employ them to analyse real-world configurations: more specifically, i) we extend the definition of several network quantities to hypergraphs, ii) compute their expected value under each null model and iii) compare it with the empirical one, in order to detect deviations from random behaviours. Differently from currently available techniques, ours is analytically tractable, scalable and effective in singling out the structural patterns of real-world hypergraphs differing significantly from those emerging as a consequence of simpler, structural constraints.
@@
Networks provide a powerful language to model interacting systems Newman (2010); Caldarelli (2010). Within a network framework, the basic unit of interaction, i.e. the edge, involves two nodes and the complexity of the structure as a whole arises from the combination of these units. Despite its many successes, network science disregards certain aspects of interacting systems, notably the possibility that more-than-two constituent units could interact at a time Lambiotte et al. (2019). Yet, it has been increasingly shown that, for a variety of systems, interactions cannot be always decomposed into a pairwise fashion and that neglecting higher-order ones can lead to an incomplete, if not misleading, representation of them Lambiotte et al. (2019); Battiston et al. (2020); Bick et al. (2023). Examples include chemical reactions involving several compounds, coordination activities within small teams of co-working people, brain activities mediated by groups of neurons. Generally speaking, thus, modelling the joint coordination of multiple entities calls for a framework that generalises the traditional, edge-centered one. An increasingly popular alternative to support a science of many-body interactions is provided by hypergraphs as these versatile, mathematical objects naturally allow nodes to interact in groups Battiston et al. (2021).
@@
Given the recent interest towards these objects, the definition of analytical tools to study them is still in its infancy Chodrow (2020); Nakajima et al. (2021); Musciotto et al. (2021); Barthelemy (2022). This paper represents our contribution to fill the gap: hereby, we extend the class of entropy-based null models Jaynes (1957); Park and Newman (2004) to hypergraphs. These models work by preserving a given set of quantities while randomising everything else, hence destroying all, possible correlations between structural properties except for those that are genuinely embodied by the constraints themselves Squartini and Garlaschelli (2011); Cimini et al. (2018); Squartini et al. (2018). The versatility of such an approach allows it to be employed either in presence of full information (to quantify the level of self-organisation of a given configuration by spotting out the patterns that are incompatible with simpler, structural constraints Squartini et al. (2013); Saracco et al. (2017); Becatti et al. (2019); Caldarelli et al. (2020); Neal et al. (2021)) or in presence of partial information (to infer the missing portion of a given configuration Parisi et al. (2020)).
@@
Our strategy for defining null models for hypergraphs is based on the randomisation of their incidence matrix. The (generally, rectangular) table contains information about the connectivity of nodes, i.e. the set of hyperedges they belong to, and the connectivity of hyperedges, i.e. the set of nodes they cluster. We will explicitly derive two members of the novel class of models hereby named Exponential Random Hypergraphs, i.e. the Random Hypergraph Model (RHM, generalising the Erdös-Rényi Model) and the Hypergraph Configuration Model (HCM, generalising the Configuration Model), and provide an analytical characterisation of their behaviour. To this aim, we will exploit the formal equivalence between the incidence matrix of a hypergraph and the biadjacency matrix of a bipartite graph Nakajima et al. (2021); Barthelemy (2022). As a next step, we will employ the HCM to assess the statistical significance of a number of patterns characterising several, real-word hypergraphs.
II II. Formalism and basic quantities
@@
A hypergraph can be defined as a pair $H(\mathcal{V},\mathcal{E}_{H})$ where $\mathcal{V}$ is the set of vertices and $\mathcal{E}_{H}$ is the set of hyperedges. Moving from the observation that the edge set $\mathcal{E}_{G}$ of a traditional, binary, undirected graph $G(\mathcal{V},\mathcal{E}_{G})$ is a subset of the power set of $\mathcal{V}$, several definitions of the hyperedge set $\mathcal{E}_{H}$ have been provided: the two most popular ones are those proposed in Berge (1967, 1973), where hyperedges ‘link’ one or more vertices, and in Voloshin (2009), where hyperedges are allowed to be empty sets as well. Hereby, we adopt the definition according to which $\mathcal{E}_{H}$ is a multiset of the power set of $\mathcal{V}$: in words, we consider non-simple hypergraphs, admitting loops and parallel edges (i.e. hyperedges involving exactly the same set nodes) of any size, including 0 (corresponding to empty hyperedges), and $|\mathcal{V}|$ (corresponding to hyperedges clustering all vertices together).
$\mathbf{I}\>=\>\begin{tabular}[]{c|c|c|c|c|c|}&$\mathbf{e_{1}}$&$\mathbf{e_{2}}$&$\mathbf{e_{3}}$&$\mathbf{e_{4}}$&$\mathbf{e_{5}}$\\ \hline\cr$\mathbf{n_{1}}$&0&1&1&0&1\\ \hline\cr$\mathbf{n_{2}}$&0&0&1&1&1\\ \hline\cr$\mathbf{n_{3}}$&0&1&1&1&0\\ \hline\cr$\mathbf{n_{4}}$&1&0&0&1&1\\ \hline\cr$\mathbf{n_{5}}$&1&1&1&1&0\\ \hline\cr$\mathbf{n_{6}}$&0&1&0&0&0\\ \hline\cr\end{tabular}$
Table 1: Example of incidence matrix describing a binary, undirected hypergraph with $6$ nodes and $5$ hyperedges. The generic entry $I_{i\alpha}$ is 1 if vertex $i$ belongs to hyperedge $\alpha$ and 00 otherwise (see also Fig. 1).
@@
As for traditional graphs, an algebraic representation of hypergraphs can be devised as well. For a formal analogy with the traditional case, let us call the cardinality of the set of nodes $|\mathcal{V}|=N$ and the cardinality of the set of hyperedges $|\mathcal{E}_{H}|=L$: then, we consider the $N\times L$ table known as incidence matrix, each row of which corresponds to a node and each column of which corresponds to a hyperedge. If we indicate the incidence matrix with $\mathbf{I}$, its generic entry $I_{i\alpha}$ will be $1$ if vertex $i$ belongs to hyperedge $\alpha$ and 00 otherwise. Notice that the number of $1$s along each row can vary between 00 and $L$, the former case indicating an isolated node and the latter one indicating a node that belongs to each hyperedge; similarly, the number of $1$s along each column can vary between 00 and $N$, the former case indicating an empty hyperedge and the latter one indicating a hyperedge that includes all nodes. As explicitly noticed elsewhere Nakajima et al. (2021); Barthelemy (2022), representing a hypergraph via its incidence matrix is equivalent at considering the bipartite graph defined by the sets $\mathcal{V}$ and $\mathcal{E}_{H}$. Table 1 represents the incidence matrix describing the binary, undirected hypergraph shown in Fig. 1.
@@
and counts the number of nodes it clusters. Both the sum of degrees and that of hyperdegrees equal the total number of $1$s, i.e. $\sum_{i=1}^{N}k_{i}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}=\sum_{\alpha=1}^{L}\sum_{i=1}^{N}I_{i\alpha}=\sum_{\alpha=1}^{L}h_{\alpha}\equiv T$. Importantly, a node degree no longer coincides with the number of its neighbours: instead, it matches the number of hyperedges it belongs to; a hyperdegree, instead, provides information on the hyperedge size.
Figure 1: Cartoon representation of the hypergraph whose incidence matrix is shown in Table 1.
III III. Binary, undirected hypergraphs randomisation
An early attempt to define randomisation models for hypergraphs can be found in Ghoshal et al. (2009). Its authors, however, have just considered hyperedges that are incident to triples of nodes - a framework that has been, later, applied to the study of the World Trade Network Nakajima et al. (2021). Considering the incidence matrix has two, clear advantages over the tensor-based representation employed in Ghoshal et al. (2009); Nakajima et al. (2021): i) generality, because the incidence matrix allows hyperedges of any size to be handled at once; ii) compactness, because the order of the tensor $\mathbf{I}$ never exceeds two, hence allowing any hypergraph to be represented as a traditional, bipartite graph.
In order to extend the rich set of null models induced by graph-specific, global and local constraints to hypergraphs, we, first, need to identify the quantities that can play this role within the novel setting. In what follows, we will consider the total number of 1s, i.e. $T$, the degree and the hyperdegree sequences, i.e. $\{k_{i}\}_{i=1}^{N}$ and $\{h_{\alpha}\}_{\alpha=1}^{L}$ either separately or in a joint fashion; moreover, we will distinguish between microcanonical and canonical randomisation techniques.
III.1 A. Homogeneous benchmarks:
@@
The model is defined by just one, global constraint, that, in our case, reads $T=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}$. Its microcanonical version extends the model by Erdös and Rényi Erdös and Rényi (1959) - also known as Random Graph Model (RGM) - to hypergraphs and prescribes to count the number of incidence matrices that are compatible with a given, total number of 1s, say $T^{*}$: they are
@@
with $V=NL$ being the total number of entries of the incidence matrix $\mathbf{I}$. Once the total number of configurations composing the microcanonical ensemble is determined, a procedure to generate them is needed: in the case of the RHM, it simply boils down to reshuffling the entries of the incidence matrix, a procedure ensuring that the total number of 1s is kept fixed while any other correlation is destroyed.
The canonical version of the RHM, instead, extends the model by Gilbert Gilbert (1959) and rests upon the constrained maximisation of Shannon entropy, i.e.
@@
where $S[P]=-\sum_{\mathbf{I}\in\mathcal{I}}P(\mathbf{I})\ln P(\mathbf{I})$, $C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining $M-1$ constraints represent proper, topological properties. The sum defining Shannon entropy runs over the set $\mathcal{I}$ of incidence matrices described in the introductory paragraph and known as canonical ensemble.
Such an optimisation procedure defines the Exponential Random Hypergraph (ERH) framework, described by the expression
@@
In the simplest case, the only, global constraint is represented by $T$ and leads to the expression
$P(\mathbf{I})$ $=\frac{e^{-\theta T(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\theta T(\mathbf{I})}}$
$=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}$
$=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}x^{I_{i\alpha}}\prod_{i=1}^{N}\prod_{\alpha=1}^{L}(1+x)^{-1}$ (6)
@@
with $e^{-\theta}\equiv x$ and $p\equiv x/(1+x)$. The canonical ensemble, now, includes all $N\times L$ rectangular matrices whose number of entries equating 1 ranges from 0 to $NL$. According to such a model, the entries of the incidence matrix are i.i.d. Bernoulli random variables, i.e. $I_{i\alpha}\sim\text{Ber}(p)$, $\forall\>i,\alpha$: as a consequence, the total number of 1s, the degrees and the hyperdegrees obey Binomial distributions - being all defined as sums of i.i.d. Bernoulli random variables. Specifically, $T\sim\text{Bin}(NL,p)$, $k_{i}\sim\text{Bin}(L,p)$, $\forall\>i$ and $h_{\alpha}\sim\text{Bin}(N,p)$, $\forall\>\alpha$, in turn, implying that $\langle T\rangle_{\text{RHM}}=NLp$, $\langle k_{i}\rangle_{\text{RHM}}=Lp$, $\forall\>i$ and $\langle h_{\alpha}\rangle_{\text{RHM}}=Np$, $\forall\>\alpha$.
@@
In order to ensure that $\langle T\rangle_{\text{RHM}}=T^{*}$, parameters have to be tuned opportunely. To this aim, the likelihood maximisation principle can be invoked Garlaschelli and Loffredo (2008): it prescribes to maximise the function $\mathcal{L}(\theta)\equiv\ln P(\mathbf{I}^{*}|\theta)$ with respect to the unknown parameter(s) that define it. Such a recipe leads us to find $p=T^{*}/NL$, with $T^{*}=T(\mathbf{I}^{*})$ indicating the empirical value of the constraint defining the RHM.
@@
The RHM (also considered in Barthelemy (2022), although a derivation from first principles is missing, there) is formally equivalent to the Bipartite Random Graph Model (BiRGM) Saracco et al. (2015). Such an identification is guaranteed by our focus on non-simple hypergraphs.
III.1.1 1. Estimation of the number of empty hyperedges
As non-simple hypergraphs admit the presence of empty as well as parallel hyperedges, let us evaluate how frequently they appear in the ensembles induced by our benchmarks.
To this aim, let us start by considering the probability for the generic hyperedge $\alpha$ to be empty or, in other terms, that its hyperdegree $h_{\alpha}$ is zero. Upon remembering that $h_{\alpha}\sim\text{Bin}(N,p)$, one finds
$p_{0}\equiv(1-p)^{N}.$ (8)
Such a quantity leads us to find the probability of observing at least one, empty hyperedge (or, in other terms, the complementary of the probability that no hyperedge is empty), i.e.
$p_{1}\equiv 1-(1-p_{0})^{L}=1-[1-(1-p)^{N}]^{L},$ (9)
and the expected number of empty hyperedges: upon remembering that they are i.i.d. Binomial random variables, one finds
$\langle N_{0}\rangle\equiv Lp_{0}=L(1-p)^{N}.$ (10)
@@
Let us, now, inspect the behaviour of the aforementioned quantities on the ensemble induced by the RHM, as the density of 1s in the incidence matrix, i.e. $p=T/NL$, varies. The two regimes of interest are the dense one, defined by $T\rightarrow NL$, and the sparse one, defined by $T\rightarrow 0$. In the dense case, one finds
$\lim_{T\to NL}p_{0}$ $=\lim_{T\to NL}\left(1-\frac{T}{NL}\right)^{N}=0,$ (11)
@@
a relationship inducing $p_{1}\xrightarrow[]{T\to NL}0$ and $\langle N_{0}\rangle\xrightarrow[]{T\to NL}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of 1s increases. Consistently, in the sparse case one finds
$\lim_{T\to 0}p_{0}=\lim_{T\to 0}\left(1-\frac{T}{NL}\right)^{N}=1,$ (12)
@@
a relationship inducing $p_{1}\xrightarrow[]{T\to 0}1$ and $\langle N_{0}\rangle\xrightarrow[]{T\to 0}L$: in words, the probability of observing empty hyperedges progressively rises as the density of 1s decreases.
@@
To evaluate the density of 1s in the incidence matrix in correspondence of which the transition from the sparse to the dense regime happens, let us consider the case $N\gg 1$: more formally, this amounts to consider the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$, i.e. $p=O(1/N)$. Since $p=T/NL$ and $h\equiv T/L$ remains finite, the probability for a generic hyperedge to be empty remains finite as well: consistently, the expected number of empty hyperedges becomes
$\lim_{N\to+\infty}\langle N_{0}\rangle=\lim_{N\to+\infty}L\left(1-\frac{h}{N}\right)^{N}=Le^{-h};$ (13)
@@
upon imposing $Le^{-h}\leq 1$, i.e. that the expected number of empty hyperedges is at most 1, one derives what may be called filling threshold, corresponding to $h_{f}^{\text{RHM}}=\ln L$. In words, a value $p>p_{f}^{\text{RHM}}\equiv h_{f}^{\text{RHM}}/N=\ln L/N$ ensures that the expected number of empty hyperedges in our random hypergraph is strictly less than one. The calculation above complements the result derived in Barthelemy (2022) about the percolation threshold for hypergraphs (see also Appendix A).
III.1.2 2. Estimation of the number of parallel hyperedges

### v3

@@ Abstract
Network theory has often disregarded many-body relationships, solely focusing on pairwise interactions: neglecting them, however, can lead to misleading representations of complex systems. Hypergraphs represent a suitable framework for describing polyadic interactions. Here, we leverage the representation of hypergraphs based on the incidence matrix for extending the entropy-based approach to higher-order structures: in analogy with the Exponential Random Graphs, we introduce the Exponential Random Hypergraphs (ERHs). After exploring the asymptotic behaviour of thresholds generalising the percolation one, we apply ERHs to study real-world data. First, we generalise key network metrics to hypergraphs; then, we compute their expected value and compare it with the empirical one, in order to detect deviations from random behaviours. Our method is analytically tractable, scalable and capable of revealing structural patterns of real-world hypergraphs that differ significantly from those emerging as a consequence of simpler constraints.
keywords:
complex networks, higher-order structures, exponential random hypergraphs
@@ Introduction
Networks provide a powerful language to model interacting systems [1, 2]. Within such a framework, the basic unit of interaction, i.e. the edge, involves two nodes and the complexity of the structure as a whole arises from the combination of these units. Despite its many successes, network science disregards certain aspects of interacting systems, notably the possibility that more-than-two constituent units could interact at a time [3]. Yet, it has been increasingly shown that, for a variety of systems, interactions cannot be always decomposed into a pairwise fashion and that neglecting higher-order ones can lead to an incomplete, if not misleading, representation of them [3, 4, 5]: examples include chemical reactions involving several compounds, coordination activities within small teams of co-working people and brain activities mediated by groups of neurons. Generally speaking, thus, modelling the joint coordination of multiple entities calls for a generalisation of the traditional edge-centred framework.
While approaches focusing on the so-called simplicial complexes have been proposed [6, 7], an increasingly popular alternative to support a science of many-body interactions is provided by hypergraphs, as these mathematical objects allow nodes to interact in groups without posing restrictions such as the ‘hierarchical’ ones characterising the former ones [8] - which, in fact, include all the subsets of a given simplex [6].
Several contributions to the definition of analytical tools for their study have already appeared [9, 10, 11, 12]: while some pertain to the purely mathematical literature and have considered probabilistic hypergraphs with the aim of studying properties such as the existence of cycles, cliques, etc. [13, 14, 15], others have adopted approaches rooted into statistical physics. Among the latter ones, some have proposed microcanonical approaches [9, 11] while others have focused on canonical ones [10, 16, 12].
@@
The present contribution aims at extending the class of entropy-based benchmarks [17, 18, 19, 20] to hypergraphs while providing a coherent framework to formally derive the canonical approaches that have been proposed so far. These models work by preserving a given set of quantities while randomising everything else, hence destroying all possible correlations between structural properties except for those that are genuinely embodied into the constraints themselves [20, 21, 22]. The versatility of such an approach allows it to be employed either in presence of full information (to quantify the level of self-organisation of a given configuration by identifying the patterns that are incompatible with simpler, structural constraints [23, 24, 25, 26, 27]) or in presence of partial information (to infer the missing portion of a given configuration [28]).
@@
Our strategy for defining null models for hypergraphs is based on the randomisation of their incidence matrix, i.e. the (generally, rectangular) table contains information about the connectivity of nodes (the set of hyperedges they belong to) and the connectivity of hyperedges (the set of nodes they cluster). We will explicitly derive two members of this novel class of models hereby named Exponential Random Hypergraphs, i.e. the Random Hypergraph Model (RHM, generalising the Erdös-Rényi Model) and the Hypergraph Configuration Model (HCM, generalising the Configuration Model), and provide an analytical characterisation of their behaviour. To this aim, we will exploit the formal equivalence between the incidence matrix of a hypergraph and the biadjacency matrix of a bipartite graph [10, 12]. Afterwards, we will employ the HCM to assess the statistical significance of a number of patterns characterising several real-word hypergraphs.
Methods
Formalism and basic quantities
@@ Methods
A hypergraph can be defined as a pair $H(\mathcal{V},\mathcal{E}_{H})$ where $\mathcal{V}$ is the set of vertices and $\mathcal{E}_{H}$ is the set of hyperedges. Moving from the observation that the edge set $\mathcal{E}_{G}$ of a traditional, binary, undirected graph $G(\mathcal{V},\mathcal{E}_{G})$ is a subset of the power set of $\mathcal{V}$, several definitions of the hyperedge set $\mathcal{E}_{H}$ have been provided: the two most popular ones are those proposed in [29, 30], where hyperedges tie one or more vertices, and in [31], where hyperedges are allowed to be empty sets as well. Hereby, we adopt the definition according to which $\mathcal{E}_{H}$ is a multiset of the power set of $\mathcal{V}$: since the concept of ‘multiset’ extends the concept of ‘set’, allowing for multiple instances of (each of) its elements, our choice implies that we are considering non-simple hypergraphs, admitting loops and parallel edges (i.e. hyperedges involving exactly the same nodes) of any size, including 0 (corresponding to empty hyperedges) and $|\mathcal{V}|$ (corresponding to hyperedges clustering all vertices together).
@@
As for traditional graphs, an algebraic representation of hypergraphs can be devised as well. In analogy with the traditional case, we call the cardinality of the set of nodes $|\mathcal{V}|\equiv N$ and the cardinality of the set of hyperedges $|\mathcal{E}_{H}|\equiv L$: then, we consider the $N\times L$ table known as incidence matrix, each row of which corresponds to a node and each column of which corresponds to a hyperedge. If we indicate the incidence matrix with $\mathbf{I}$, its generic entry $I_{i\alpha}$ will be $1$ if vertex $i$ belongs to hyperedge $\alpha$ and 00 otherwise. Notice that the number of $1$s along each row can vary between 00 and $L$, the former case indicating an isolated node and the latter one indicating a node that belongs to each hyperedge; similarly, the number of $1$s along each column can vary between 00 and $N$, the former case indicating an empty hyperedge and the latter one indicating a hyperedge that includes all nodes. As explicitly noticed elsewhere [9, 10, 12], representing a hypergraph via its incidence matrix is equivalent to considering the bipartite graph defined by the sets $\mathcal{V}$ and $\mathcal{E}_{H}$ more formally, the function that assigns each hypergraph to the bipartite graph associated with it is a bijection when both nodes and hyperedges are uniquely labeled [9]. For instance, the incidence matrix $\mathbf{I}$ describing the binary, undirected hypergraph shown in Fig. 1 is the following:
Figure 1: Cartoon representation of a hypergraph. Black dots represent nodes, while the colored shapes represent hyperedges. The incidence matrix describing the present hypergraph is defined by Eq. 1.
[…]
@@
and counts the number of nodes it clusters. Both the sum of degrees and that of hyperdegrees equal the total number of $1$s, i.e. $\sum_{i=1}^{N}k_{i}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}=\sum_{\alpha=1}^{L}\sum_{i=1}^{N}I_{i\alpha}=\sum_{\alpha=1}^{L}h_{\alpha}\equiv T$. Importantly, a node degree no longer coincides with the number of its neighbours: instead, it matches the number of hyperedges it belongs to; a hyperdegree, instead, provides information about the hyperedge size. Analogously, $T$ paves the way for the alternative definition of ‘density of connections’ reading $\rho=T/NL\equiv h/N$, i.e. the ratio between the (average) number of nodes each hyperedge clusters and the total number of nodes.
Binary, undirected hypergraphs randomisation
An early attempt to define randomisation algorithms for hypergraphs within a framework closely resembling ours can be found in [32]. Its authors, however, have just considered hyperedges that are incident to triples of nodes - a framework that has been, later, applied to the study of the World Trade Network [10].
Considering the incidence matrix has two clear advantages over the tensor-based representation employed in [32, 10]: i) generality, because the incidence matrix allows hyperedges of any size to be handled at once; ii) compactness, because the order of the tensor $\mathbf{I}$ never exceeds two, hence allowing any hypergraph to be represented as a traditional, bipartite graph.
In order to extend the rich set of null models induced by graph-specific global and local constraints to hypergraphs, we, first, need to identify the quantities that can play this role within the novel setting. In what follows, we will consider the total number of $1$s, i.e. $T$, the degree and the hyperdegree sequences, i.e. $\{k_{i}\}_{i=1}^{N}$ and $\{h_{\alpha}\}_{\alpha=1}^{L}$ either separately or in a joint fashion; moreover, we will distinguish between microcanonical and canonical randomisation techniques.
Homogeneous benchmarks:
@@
Microcanonical formulation
The model is defined by just one, global constraint, that, in our case, reads
$T=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha};$ (4)
its microcanonical version extends the model by Erdös and Rényi [33] - also known as Random Graph Model (RGM) - to hypergraphs and prescribes to count the number of incidence matrices that are compatible with a given, total number of $1$s, say $T^{*}$: they are
@@
with $V\equiv NL$ being the total number of entries of the incidence matrix $\mathbf{I}$. Once the total number of configurations composing the microcanonical ensemble has been determined, a procedure to generate them is needed: in the case of the RHM, it simply boils down to reshuffling the entries of the incidence matrix, a procedure ensuring that the total number of $1$s is kept fixed while any, other correlation is destroyed.
Canonical formulation
The canonical version of the RHM, instead, extends the model by Gilbert [34] and rests upon the constrained maximisation of Shannon entropy, i.e.
@@
where $S[P]=-\sum_{\mathbf{I}\in\mathcal{I}}P(\mathbf{I})\ln P(\mathbf{I})$, $C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining $M-1$ constraints represent proper topological properties. The sum defining Shannon entropy runs over the set $\mathcal{I}$ of incidence matrices described in the introductory paragraph and known as canonical ensemble. Such an optimisation procedure defines the Exponential Random Hypergraph (ERH) framework, described by the expression
@@
In the simplest case, the only global constraint is represented by $T$ and leads to the expression
$P(\mathbf{I})$ $=\frac{e^{-\theta T(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\theta T(\mathbf{I})}}=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}x^{I_{i\alpha}}\prod_{i=1}^{N}\prod_{\alpha=1}^{L}(1+x)^{-1}$ (8)
@@
with $e^{-\theta}\equiv x$ and $p\equiv x/(1+x)$. The canonical ensemble, now, includes all $N\times L$, rectangular matrices whose number of entries equal to 1 ranges from 0 to $NL$. According to such a model, the entries of the incidence matrix are i.i.d. Bernoulli random variables, i.e. $I_{i\alpha}\sim\text{Ber}(p)$, $\forall\>i,\alpha$; as a consequence, the total number of $1$s, the degrees and the hyperdegrees obey Binomial distributions, being all defined as sums of i.i.d. Bernoulli random variables: specifically, $T\sim\text{Bin}(NL,p)$, $k_{i}\sim\text{Bin}(L,p)$, $\forall\>i$ and $h_{\alpha}\sim\text{Bin}(N,p)$, $\forall\>\alpha$, in turn, implying that $\langle T\rangle_{\text{RHM}}=NLp$, $\langle k_{i}\rangle_{\text{RHM}}=Lp$, $\forall\>i$ and $\langle h_{\alpha}\rangle_{\text{RHM}}=Np$, $\forall\>\alpha$.
Parameter estimation
@@
In order to ensure that $\langle T\rangle_{\text{RHM}}=T^{*}$, parameters have to be tuned opportunely. To this aim, the likelihood maximisation principle can be invoked [19]: it prescribes to maximise the function $\mathcal{L}(\theta)\equiv\ln P(\mathbf{I}^{*}|\theta)$ with respect to the unknown parameter that defines it. Such a recipe leads us to find
$p=\rho^{*}=\frac{T^{*}}{NL}$ (10)
with $T^{*}=T(\mathbf{I}^{*})$ indicating the empirical value of the constraint defining the RHM.
@@
The RHM (also considered in [12], although without providing any derivation from first principles, and in [16], although without providing any recipe for the estimation of its parameter) is formally equivalent to the Bipartite Random Graph Model (BiRGM) [35]. Such an identification is guaranteed by our focus on non-simple hypergraphs.
Estimation of the number of empty hyperedges
Non-simple hypergraphs admit the presence of empty as well as parallel hyperedges. As this type of structures is associated with configurations that may be regarded as problematic (since not observed in empirical data), we evaluate how frequently they appear in the ensembles induced by our benchmarks. Let us denote the number of empty hyperedges, i.e. the number of hyperedges whose hyperdegree equals zero, with $N_{\emptyset}$: since the hyperdegrees are i.i.d. Binomial random variables, $N_{\emptyset}\sim\text{Bin}(L,p_{\emptyset})$ where
${\color[rgb]{0,0,0}p_{\emptyset}\equiv(1-p)^{N}}$ (11)
is the probability for the generic hyperedge to be empty or, equivalently, for its hyperdegree to equal zero; the expected number of empty hyperedges reads
${\color[rgb]{0,0,0}\langle N_{\emptyset}\rangle=Lp_{\emptyset}=L(1-p)^{N}.}$ (12)
@@
Let us, now, inspect the behaviour of $p_{\emptyset}$ on the ensemble induced by the RHM as the density of $1$s in the incidence matrix, i.e. $p=T/NL$, varies. The two regimes of interest are the dense one, defined by $T\rightarrow NL$, and the sparse one, defined by $T\rightarrow 0$. In the dense case, one finds
$\lim_{T\to NL}p_{\emptyset}$ $=\lim_{T\to NL}\left(1-\frac{T}{NL}\right)^{N}=0,$ (13)
@@
a relationship inducing $\langle N_{\emptyset}\rangle\xrightarrow[]{T\to NL}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of $1$s increases. Consistently, in the sparse case one finds
$\lim_{T\to 0}p_{\emptyset}=\lim_{T\to 0}\left(1-\frac{T}{NL}\right)^{N}=1,$ (14)
@@
a relationship inducing $\langle N_{\emptyset}\rangle\xrightarrow[]{T\to 0}L$: in words, the probability of observing empty hyperedges progressively rises as the density of $1$s decreases.
@@
To evaluate the density of $1$s in the incidence matrix in correspondence of which the transition from the sparse to the dense regime happens, let us consider the case $N\gg 1$: more formally, this amounts to consider the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$ equivalently, defined by posing $p=O(1/N)$. Since $p=T/NL$ and $h=T/L$ remains finite, the probability for the generic hyperedge to be empty obeys the relationship
${\color[rgb]{0,0,0}\lim_{N\to+\infty}p_{\emptyset}=\lim_{N\to+\infty}\left(1-\frac{h}{N}\right)^{N}=e^{-h},}$ (15)
i.e. remains finite as well: consistently, the expected number of empty hyperedges becomes
${\color[rgb]{0,0,0}\lim_{N\to+\infty}\langle N_{\emptyset}\rangle=\lim_{N\to+\infty}Lp_{\emptyset}=Le^{-h};}$ (16)
@@
upon imposing $Le^{-h}\leq 1$, i.e. that the expected number of empty hyperedges is at most 1, one derives what may be called filling threshold, corresponding to $h_{f}^{\text{RHM}}\equiv\ln L$. In words, a value
${\color[rgb]{0,0,0}p>p_{f}^{\text{RHM}}=\frac{h_{f}^{\text{RHM}}}{N}=\frac{\ln L}{N}}$ (17)
ensures that the expected number of empty hyperedges in our random hypergraph is strictly less than one. As a last observation, let us notice that evaluating $p_{\emptyset}$ in correspondence of the filling threshold returns the value $1/L$.
A comparison with simple graphs: estimating the number of isolated nodes
A similar line of reasoning can be repeated for traditional graphs, the aim being, now, that of estimating $N_{0}$, i.e. the number of isolated nodes. To this aim, let us consider the asymptotic framework defined by letting $N\to+\infty$ while posing $L=O(N)$ equivalently, defined by posing $q=O(1/N)$. Since $q=2L/N(N-1)$ and $k\equiv 2L/(N-1)$ remains finite, the probability for the generic node $i$ to be isolated obeys the relationship
${\color[rgb]{0,0,0}\lim_{N\to+\infty}q_{0}=\lim_{N\to+\infty}(1-q)^{N-1}=\lim_{N\to+\infty}\left(1-\frac{k}{N-1}\right)^{N-1}=e^{-k},}$ (18)
i.e. remains finite as well: this, in turn, implies that the expected number of isolated nodes $\langle N_{0}\rangle=Nq_{0}$ obeys the relationship
${\color[rgb]{0,0,0}\lim_{N\to+\infty}\langle N_{0}\rangle=\lim_{N\to+\infty}Nq_{0}=Ne^{-k};}$ (19)
upon imposing $Ne^{-k}\leq 1$, i.e. that the expected number of isolated nodes is at most 1, one derives the connectivity threshold, corresponding to $k_{c}^{\text{RGM}}\equiv\ln N$. In words, a value
${\color[rgb]{0,0,0}q>q_{c}^{\text{RGM}}=\frac{k_{c}^{\text{RGM}}}{N}=\frac{\ln N}{N}}$ (20)
ensures that the expected number of isolated nodes in our random graph is strictly less than one [1]. As a last observation, let us notice that evaluating $q_{0}$ in correspondence of the connectivity threshold returns the value $1/N$.
We also note that $N$ is the only quantity playing a relevant role in the case of graphs, while an interplay between $L$ and $N$ can be observed in the case of hypergraphs: in both cases, however, a condition on connectivity is present, driven by the request that the objects under investigations (nodes on the one hand and hyperedges on the other) have a non-zero number of connections.
Estimation of the number of parallel hyperedges